1)    The theoretical maximum yield of MgCl2 that could be formed from 5.00g Mg in the reaction,

Mg (s)+2HCl (aq)→MgCl2 (aq)+H2 (g)

is 19.6g. In a laboratory experiment, a student, was able to make 14.2g. What was the percentage yield?






2). A pressure of 101,325 Nm-2 is equivalent to 1.00 atmospheres. Convert a pressure of 0.65 atmospheres to newton per metre squared (Nm-2), giving your answer to 3 significant figures.










3). The relationship between concentration, amount of material (number of moles) and volume is,

Concentration (mol dm-3)   =  moles (mol) / volume (dm3)

A solution contains 1.5 x 10-2 mol of sodium hydroxide (NaOH) dissolved in 75.0 cm3 of water. What is the concentration of the sodium hydroxide solution?

 







4). The speed of light in a vacuum is 299,792,500 ms-1. Convert to standard form and give the value to 3 significant figures.












 








1
Expert's answer
2021-06-21T05:28:29-0400

% yield = Actual yield/Theoretical yield × 100

Theoretical = 19.6-5 = 14.6

Actual = 14.2

% Yield = 14.2/14.6 × 100

= 97.2%


2. 1 = 101,325

0.65 = ?

= 0.65 × 101325

= 65861.25Nm-2


3). Concentration = Moles of solute /Littres of the solution

= 0.015

Molar Mass of NaOH = 39.997

= 0.015 × 39.997 = 0.6

= 0.0006+0.075

= 0.07506

= 0.015/0.07506

= 0.1998 mol/Littre


4. 299792500

= 3 × 108 MS-1